# Addressing Blind Guessing: Calibration of Selection Bias in Multiple-Choice Question Answering by Video Language Models

### v1

@@ Abstract
Evaluating Video Language Models (VLMs) is a challenging task. Due to its transparency, Multiple-Choice Question Answering (MCQA) is widely used to measure the performance of these models through accuracy. However, existing MCQA benchmarks fail to capture the full reasoning capabilities of VLMs due to selection bias, when models disproportionately favor certain answer options based on positional patterns observed during training. In this work, we conduct a comprehensive empirical analysis of several VLM architectures across major datasets designed to assess complex video-focused reasoning. We identify where the bias is most pronounced and demonstrate to what extent model responses reflect genuine understanding of video content and related questions, as opposed to reliance on arbitrary patterns or superficial cues, such as answer position. By decomposing the MCQA task and adapting fairness bias metrics to VLMs, we introduce a post-processing calibration technique BOLD to balance this bias. Our results show that reducing selection bias improves not only debiasing metrics but also overall model performance, including Accuracy and F1 Mean score. Our method, by suppressing "blind guessing", offers a more cost- and time-effective approach to mitigating selection bias compared to existing techniques. This study represents the first focused investigation of selection bias in video-to-text LLM-powered models.
@@
Olga Loginova University of Trento olga.loginova@unitn.it                        Oleksandr Bezrukov Gran Sasso Science Institute oleksandr.bezrukov@gssi.it                        Alexey Kravets University of Bath ak3095@bath.ac.uk
@@ 1 Introduction
Multiple-choice question answering (MCQA) is a convenient method to assess language models’ performance thanks to its simplicity and transparency. Each MCQA instance presents a question with several answer options, and the task is to select the correct one. However, large language models (LLMs) often exhibit selection bias, favoring certain answer positions Zheng et al. (2024a); Wang et al. (2024a). This bias undermines the fairness, reliability, and reasoning capabilities of the models.
While LLMs process text, video-language models (VLMs) handle sequences of frames to capture both spatial and temporal features. Recent advances align pre-training LLMs with video encoders to enhance video understanding Bordes et al. (2024). Although VLMs inherit much from LLMs, they face unique challenges, including higher computational costs and the need for effective spatial and temporal mapping for dynamic inputs.
Despite their backbone shared with LLMs, VLMs’ specific characteristics require separate scrutiny of selection bias. Notably, no research has yet focused on bias in video-language models.
Figure 1: Decomposition approach: we separate the three key components of MCQA, reorganize them into pairs by removing one at a time, and apply probability debiasing to the aggregated data.
We address this gap by proposing the first study of MCQA selection bias in video datasets, focusing on temporal reasoning, spatial object tracking, and camera behavior rather than general plot understanding. This approach highlights the specific challenges of VLMs.
Traditionally, selection bias is mitigated through shuffling Zheng et al. (2024a); Liu et al. (2023), where answer options are permutated multiple times to yield more impartial results. In contrast, we propose treating it similarly to the fairness bias, where models exhibit preference toward specific groups, such as gender or race. We adapt fairness bias metrics for VLMs via the decomposition approach showed in Figure  1.
To evaluate our approach, we select three VLMs with different architectures and apply post-processing bias calibration techniques to four datasets. Our method, which adjusts predictions without retraining, improves interpretability and fairness with reduced computational costs. Even though there is no direct correlation between debiasing and performance, our approach generally improves both bias metrics and accuracy, as well as the F1 mean score, signaling improved recall.
@@
  • A comprehensive study of selection bias in VLMs using 11 dataset modifications focused on video-specific challenges.
  • Adaptation of debiasing metrics to assess VLMs’ behavior on video MCQA benchmarks.
  • An efficient calibration technique that improves the reliability of current MCQA datasets and the VLM performance assessment.
Our work aims to reveal deeper reasoning issues in contemporary VLMs through refined MCQA benchmarks.
@@ 3.1 Datasets
We selected four video MCQA datasets that challenge VLMs’ temporal and spatial reasoning abilities. These datasets represent everyday situations and span videos from a few seconds to over an hour, most of which last up to one minute.
Table 1: Summary of Question-Answer pairs, Videos, and Number of Options for each dataset.
@@
NExT-QA and NExT-GQA
NExT-QA Xiao et al. (2021) is one of the first datasets designed to assess a VLM’s ability to capture causal, temporal, and descriptive topics through what/when/why/how and before/after questions. We evaluate the models on the test split.
@@
NExT-GQA is a subset of the NExT-QA test split annotated with timestamp to locate the answer moment in the video.
STAR
The question-answer (QA) pairs in STAR Wu and Yu (2021) are generated using functional programs based on situation hyper-graphs. They cover four categories of situational reasoning: Interaction, Sequence, Prediction, and Feasibility. We use the publicly available validation split.
Perception Test
The Perception Test dataset Puatruaucean et al. (2023) tests the memory, abstraction, physics, and semantics reasoning abilities of the models. We use a 40% split of the correct-answer-annotated validation set to save the resources. Sampling is stratified to maintain the original balance of question types and the correct answer distribution.
Video-MME
Video-MME Fu et al. (2024) is a comprehensive MCQA evaluation benchmark with key distinctions: (a) video duration range from 11 seconds to 1 hour, (b) it covers 6 primary visual domains with 30 subfields, (c) includes subtitles, and (d) features meticulously crafted annotations. The released test split has an imbalanced option distribution, so we equalized it to properly assess option bias.
The amount of QA pairs, videos and answer options of each dataset is shown in Table 1. Additional details on the datasets including QA examples are given in Appendix  A.
@@ 3.2 Models
Three models were selected to represent different approaches to uniting video and language modalities Bordes et al. (2024). Despite architectural differences, SeViLA, Video-LLaMA, and Video-LLaVA share key characteristics: they sample frames to extract visual features, either by keyframe selection or uniform sampling, and align visual features with textual representations using pre-trained LLMs. Each model uses pre-trained visual encoders and fine-tunes pre-trained components with video-text datasets.
Video-LLaMA
Video-LLaMA Zhang et al. (2023a) aligns a BLIP-2-based visual encoder Li et al. (2023) with LLaMA Touvron et al. (2023). It incorporates audio signals using Imagebind Girdhar et al. (2023). Both the video and audio Q-formers are trained on the Webvid-2M dataset Bain et al. (2021) to align the modalities. LLaMA is fine-tuned with visual instructional data.
Video-LLaVA
Video-LLaVA (Lin et al., 2023a) aligns video and LLMs in a shared representation. It uses the pre-trained CLIP ViT-L/14 vision encoder Radford et al. (2021) and the Vicuna language model Chiang et al. (2023), improving multimodal instruction following.
SeViLA
SeViLA Yu et al. (2023) adopts a self-chaining approach with BLIP-2 Li et al. (2023) for keyframe localization and question answering. It includes a Localizer, which uniformly samples 32 frames and selects the 4 most relevant key ones, and an Answerer that generates responses based on those frames. The Localizer and Answerer both have Flan-T5  Chung et al. (2022) for LLM.
We provide the Accuracy and F1 Mean Score results of the zero-shot inference on the aforementioned datasets in Table 2. Further details on the model settings and prompts can be found in Appendix B.
@@ 3.3 Experimental Settings
Eleven dataset modifications are used to examine the role of each component in MCQA benchmarks: video, question, and answer options.
@@
  • Correct Frames: Only frames containing the answer are provided. This tests if reducing extraneous visual information improves performance. The modification applies to NExT-QA (NExT-GQA) and STAR due to the availability of timestamp annotations.
  • Empty Frames: Black frames are passed instead of video input. This tests the model’s response to the absence of visual information within the multimodal setting.
@@
  • Rephrased Questions: Llama3 Dubey et al. (2024) rephrases the original questions. Five variations are generated, and one is randomly selected. [^1] A manual check of 10% of each rephrased dataset confirmed that meaning was retained, with minimal impact on accuracy due to meaning loss. Notably, very few hallucinations were spotted in the NExT-QA rephrased copy.
  • Empty Questions: Following Balepur et al. (2024), we replace the question with an empty string to test if models can infer the question from answer choices and video. [^2] The previous research for LLMs has shown surprisingly high accuracy in this scenario. We investigate if the video modality contributes to even more meaningful predictions.
@@
examine the role of answer:
  • Answer Shuffling: Options are randomly shuffled to test robustness against bias through changing the option order.
  • Correct Answer in Each Option: The correct answer is placed in a fixed position; its content is swapped with another option.
  • Correct Answer with Shuffling: The correct answer is fixed, and the remaining options are shuffled.
  • Additional Empty Option: A new empty option is added in the end to check for overfitting to specific options.
  • All Identical Answers: All options are identical testing for deviation from uniform distribution.
  • All Correct Answers: Every option is correct testing for deviations in a setting where any answer is valid.
  • Empty Answers: All options are empty to check the model’s ability to choose a non-existent answer.
@@
Modifications can also be categorized as accuracy-relevant (e.g., correct frames, shuffling, or rephrased questions) and accuracy-irrelevant (e.g., identical answers or empty parts). Further details, including accuracy results and confusion matrices, are provided in the Appendix C.
@@ 3.4 Observations
Figures  2– 4 display the answer distributions of each model across options and datasets. High bars concentrated in a specific option indicate bias, while consistently low bars suggest the model disregards that option.
Across all models and datasets, shuffling answer options – either by shifting the correct option or random permutation, results in similar option distributions. This indicates that models rely heavily on answer position rather than content. However, when the correct answer is moved to less favored positions, those positions show increased selection rates, as seen in peaks for Correct ai and Correct ai Shuffled. Additionally, rephrased questions yield distributions similar to the default, regardless of correctness.
Using all correct frames did not significantly alter the distributions, suggesting that the models’ localization capabilities are adequate. Remarkably, in scenarios where all answer options are identical (accuracy-irrelevant), the selection distribution is not uniform.
Video-LLaMA
Video-LLaMA exposes a significant bias toward option a1 (Figure 2): when the correct answer is in this position, accuracy exceeds 70% on NExT-QA and Perception Test. This bias is more evident across STAR, Perception Test, and Video-MME, especially with empty frames.
The model tends to ignore options beyond the third one. In Perception Test, which has only three choices, bias against the third option is already noticeable.
Notably, with empty answers, option a2 receives an unexpectedly high number of selections, even surpassing a1 in Video-MME.
Video-LLaVA
Video-LLaVA’s performance (Figure 3) is similar to Video-LLaMA’s but shows an even stronger bias toward option a1, with higher-numbered options being largely ignored. Unlike Video-LLaMA, Video-LLaVA overwhelmingly favors a1 when answers are empty. When questions are empty, option a0 starts competing with a1, especially in the Video-MME dataset. In NeXT-QA, moving the correct answer among options leads to a preference for the option containing it.
SeViLA
SeViLA is unique in becoming distracted when an empty option is added. This suggests SeViLA may have been trained on answer distributions with a fixed number of options. It shows reduced accuracy for all correct frames in STAR (see Table  13), echoing findings by Goyal et al. (2016) and Lin et al. (2023b), where "blind" language-only models could answer image questions.
Nevertheless, SeViLA demonstrates the greatest robustness on NExT-QA and STAR. Its unstable option distributions occur only when all correct answers are identical (preferring a0) and with empty answers (preferring the last option). Figure 4 shows slightly higher saturation for these options across all datasets.
@@ 4 Methodology
The bias calibration is based on the following requirements to the MSQA benchmark:
  • Balanced options: All answer options are presented equally without any inherent bias (e.g., one answer option is not more likely than others just by how it is presented).
  • Independent options: Each option for a particular question is independent of others (e.g., there are no inherent relationships between them that would influence selection).
  • Equal difficulty: All questions are equally difficult, and no question skews results toward a particular option group.
  • No logically better or worse options except the correct one: This ensures that other options are plausible distractors, but only one option is objectively correct.

### v2

@@ Abstract
Evaluating Video Language Models (VLMs) is a challenging task. Due to its transparency, Multiple-Choice Question Answering (MCQA) is widely used to measure the performance of these models through accuracy. However, existing MCQA benchmarks fail to capture the full reasoning capabilities of VLMs due to selection bias, when models disproportionately favor certain answer options based on positional patterns observed during training. In this work, we conduct a comprehensive empirical analysis of several VLM architectures across major datasets designed to assess complex video-focused reasoning. We identify where the bias is most pronounced and demonstrate to what extent model responses reflect genuine understanding of video content and related questions, as opposed to reliance on arbitrary patterns or superficial cues, such as answer position. By decomposing the MCQA task and adapting fairness bias metrics to VLMs, we introduce a post-processing calibration technique BOLD to balance this bias. Our results show that reducing selection bias improves not only debiasing metrics but also overall model performance, including Accuracy and F1 Mean score. Our method, by suppressing "blind guessing", offers a more cost- and time-effective approach to mitigating selection bias compared to existing techniques. This study represents the first focused investigation of selection bias in video-to-text LLM-powered models. [^1] The codes for the settings, models and calibration technique are avaialable on Github: https://github.com/ologin/BOLD
@@
Olga Loginova University of Trento olga.loginova@unitn.it                        Oleksandr Bezrukov Gran Sasso Science Institute oleksandr.bezrukov@gssi.it
Ravi Shekhar University of Essex r.shekhar@essex.ac.uk                        Alexey Kravets University of Bath ak3095@bath.ac.uk
@@ 1 Introduction
Multiple-choice question answering (MCQA) is a transparent and convenient method to assess language models’ performance. Each MCQA instance presents a question and several answer options, requiring the selection of the correct one. However, large language models (LLMs) often exhibit selection bias, i.e., a tendency to favor certain answer positions irrespective of content Zheng et al. (2024a); Wang et al. (2024a). Such bias compromises fairness, reliability, and the true reasoning capabilities of these models. Although selection bias in LLMs has received increasing attention, it remains largely unexplored for video-language models (VLMs).
VLMs integrate textual and visual inputs by processing sequences of frames and capturing both spatial and temporal features. Recent advances align pre-trained LLMs with video encoders Bordes et al. (2024). Although VLMs inherit many properties from LLMs, their reliance on complex visual inputs introduces unique challenges, such as higher computational overhead and the need for robust spatiotemporal reasoning. The growing role of VLM-based MCQA tasks thus calls for a tailored investigation of selection bias in VLMs.
We address this gap with the first study of video-MCQA selection bias, focusing on video-specific aspects such as dynamic unfolding of events, spatial object tracking, and camera behavior, rather than general plot understanding.
Figure 1: Decomposition approach in 3 steps: key-component decomposition, rearrangement into pairs, applying probability debiasing to the aggregated data.
Through the empirical analysis, we locate where the model’s predictions are skewed by inherent option preferences. We find that, while each model demonstrates specific biases, accuracy-relevant settings yield option distributions similar to those in the default setting. This suggests that the models rely heavily on answer position, regardless of content. Conversely, in accuracy-irrelevant settings, the predicted distributions deviate significantly from uniformity, and this is where the bias is most apparent. Subsequently, we adapt the metrics from fairness bias studies (e.g., regarding race or age) to video-MCQA and deliver a calibration technique.
Figure 1 illustrates our decomposition strategy. We find that depriving the MCQA task of its key constituent components (video, question, answer options) reveals latent bias factors. We systematically isolate the three components: we remove one component at a time and rearrange the remaining pairs, obtaining three “ill-defined” variants of the original data. These decompositions allow us to identify the underlying bias patterns across different sub-spaces of the task. We then aggregate these insights and apply a probability-based debiasing procedure that adjusts the final predictions.
By experimenting with three VLM architectures and four video-specific MCQA datasets, we show that our post-training practical solution enhances interpretability, reduces computational complexity [^2] Traditionally, selection bias is mitigated through shuffling Zheng et al. (2024a); Liu et al. (2023), where answer options are permuted multiple times to produce more impartial results. Our method requires far fewer resources compared to all the exhaustive answer-option permutations., and improves bias metrics. Intriguingly, while our method is designed to mitigate bias, it generally improves performance with better accuracy and F1 Mean scores.
@@
  • A thorough investigation of selection bias in VLM-based MCQA with 11 dataset modifications to target video-specific reasoning challenges.
  • Adaptation of fairness metrics to assess selection bias in VLMs on video MCQA benchmarks.
  • A novel decomposition post-processing calibration technique that leverages the rearrangement of MCQA, mitigates selection bias and improves performance.
Through the refined MCQA benchmarks, we uncover deeper reasoning issues in contemporary VLMs. Our work helps move toward more robust, fair, reliable and interpretable VLMs by bridging a gap in the study of selection bias in video understanding models and offering a cost-effective solution.
@@ 3.1 Datasets
Table 1: Summary of Question-Answer (QA) pairs, Videos, and Number of Options for each dataset.
@@
Our goal is to investigate VLMs’ bais under different conditions. To this end, we chose datasets specifically to cover a wide variety of question types, including what, when, why, how, and yes/no questions, that represent a broad spectrum of video-specific reasoning categories such as causal, temporal, descriptive, situational reasoning, as well as memory, abstraction, physics, and semantics.
@@
We selected the following four video MCQA datasets: NExT-QA Xiao et al. (2021) (and its subset NExT-GQA Xiao et al. (2023) with timestamp annotations to locate the answer moment), STAR Wu and Yu (2021), Perception Test Puatruaucean et al. (2023), and Video-MME Fu et al. (2024). These datasets consist of videos of varying lengths, from a few seconds to over an hour. The questions are either manually crafted or generated via functional programs, and the video content range from short everyday activities to animations.
The datasets also differ in the number of answer options per question. Table 1 gives the overview of the number of QA pairs, videos, and answer options for each dataset. Further details on each dataset, including QA examples, are provided in Appendix A.
For experimenting, we used the test splits of NExT-QA/NExT-GQA and Video-MME and the publicly available validation splits with correct answer annotations of STAR and Perception Test.
@@ 3.2 Models
To explore selection bias in a broader context, we chose to examine it with respect to diverse model architectures. Despite the differences, the selected models share key characteristics: they sample frames to extract visual features, either by keyframe selection or uniform sampling, and align visual features with textual representations using pre-trained LLMs. Each model leverages pre-trained visual encoders and fine-tunes pre-trained components with video-text datasets. The models differ substantially in methods of integrating video and language modalities Bordes et al. (2024):
Figure 2: Video-LLaMA option distribution for accuracy-irrelevant settings across tested datasets. All Each represents the aggregated distribution under the All Identical Answers setting.
Video-LLaMA Zhang et al. (2023a) has an additional audio channel and uses three query transformers (Q-formers) trained to align the video, audio, and language modalities.
Video-LLaVA (Lin et al., 2023a) aligns videos and LLMs within a shared representation space and has improved multimodal instruction-following capabilities.
SeViLA Yu et al. (2023), specifically designed for question answering, implements a self-chaining approach with Localizer and Answerer: the first selects relevant frames, which the second uses to generate or choose answers.
Table 2 presents the zero-shot inference results of each model on the chosen datasets. Further details on the models, their settings, and prompts can be found in Appendix B.
@@ 3.3 Experimental Settings
Eleven dataset modifications are used to examine the role of each component in MCQA benchmarks: video, question, and answer options. The outmost purpose of the modifications is to identify 1) if bias is universal across models and datasets, 2) in what option(s) bias is, 3) how much each component contribute to the bias, and 4) in what settings it is the strongest with respect to each video-MCQA component.
@@
1. Correct Frames: Only frames containing the answer are given as video input. This tests if reducing extraneous visual information improves performance. The modification applies only to NExT-QA (NExT-GQA) and STAR due to the availability of timestamp annotations.
2. Empty Frames: Black frames are passed instead of meaningful video input. This tests the model’s response to the absence of visual information within the multimodal setting.
@@
3. Rephrased Questions: Llama3 Dubey et al. (2024) rephrases the original questions five times, then one rephrasing is randomly selected. [^3] A manual check of 10% of each rephrased dataset confirmed that meaning was retained, with minimal impact on accuracy due to meaning loss. Notably, very few hallucinations were spotted in the NExT-QA rephrased copy. By changing the wording, we check if the selection bias is linked to the token bias.
4. Empty Questions: Following Balepur et al. (2024a), we replace the question with an empty string to test if models can infer the question from answer choices and video. [^4] The previous research for LLMs Balepur et al. (2024b) has shown surprisingly high accuracy in this scenario. We investigate if the video modality contributes to even more meaningful predictions.
@@
examine the role of answer and its position in the emergence of bias:
5. Answer Shuffling: Options are randomly shuffled once. Previous experiments Zheng et al. (2023); Zong et al. (2023) showed that the models are not robust against changing the option order. Regardless of accuracy, we check if the distribution of answers in each option is the same, which reveals the bias consistency.
6. Correct Answer in Each Option: The correct answer is placed in a fixed position; its content is swapped with another option. By changing the position of correct answer we test to what extend each option contribute to the models’ choice.
7. Correct Answer with Shuffling: The correct answer is fixed, and the remaining options are shuffled. This fusion of the previous two approaches tests how stable a particular option choice is given that it is correct.
8. Additional Empty Option: A new empty option is added in the end to check for overfitting to specific options. If the added blank line is predicted, this indicates a biased distribution of predictions among a certain number of options.
9. All Identical Answers: All options are identical testing for deviation from uniform distribution. When the answer from each option becomes the answer of all options, the answer bias with respect to tokens is eliminated. Moreover, since there is no correct answer in the case of distractors, we can check the answer choice that happens without meaningful reasoning.
10. All Correct Answers: Every option is correct. In addition to (9), where any answer is equal, in this setting any answer is valid. The expectation is to get the consistent uniform distribution.
11. Empty Answers: Each answer option is an empty string. With a non-existent answer, the option name and its position in the question are the only sources for choosing an answer.
@@
Modifications can also be categorized as accuracy-relevant (e.g., correct frames, shuffling, or rephrased questions) and accuracy-irrelevant (e.g., identical answers or empty parts). Further details, including accuracy results, confusion matrices and examples for each modification are provided in the Appendix C.
Figure 3: Video-LLaVA option distribution for accuracy-irrelevant settings across tested datasets. All Each represents the aggregated distribution under the All Identical Answers setting.
@@ 3.4 Observations
Accuracy-relevant settings (Correct Frames, Rephrased Questions, and Answer Shuffling) vary in accuracy but yield almost the same option distributions as the default setting. This suggests that models rely more on answer position than on visual or textual content.
When the correct answer is placed in each option (Correct Answer in Each Option) or combined with shuffling (Correct Answer with Shuffling), the selection rate for that option increases, which indicates adequate reasoning. Moving the correct answer into less favored positions increases their selection proportion, and adding shuffling to it has minimal additional effect.
In contrast, accuracy-irrelevant settings clearly expose bias. Under identical options (All Identical Answers and All Correct Answers) or incomplete conditions (Empty Frames, Empty Questions, Empty Answers), the predicted distributions deviate significantly from uniformity. Figures 2-4 highlight the patterns of the accuracy-agnostic settings as they are indicative for the study of bias, while full performance details are in Appendix C.
In case of All Identical Answers each-option answer showed consistently a very close distribution, so in the figures the result of All Each is obtained by averaging the outcomes across every option experiments within this setting. Tall bars concentrated in certain options indicate bias, while persistently low bars reveal neglected options. Each model exhibits pronounced biases toward specific options.
Video-LLaMA (Figure 2) strongly favors a1 and tends to ignore options beyond the third one. In the Perception Test (three options), it already shows bias against the third option. With identical answers, a0 sometimes competes with a1, and when answers are empty, a2 surprisingly surpasses a1 in Video-MME. Empty settings cover the behavior of the bias to the greatest extent.
Video-LLaVA (Figure 3) behaves similarly but exhibits even stronger bias toward a1, largely ignoring higher-numbered options. Unlike Video-LLaMA, it overwhelmingly favors a1 when answers are empty. With empty questions, a0 begins to compete with a1, especially in Video-MME.
Figure 4: SeViLA option distribution for accuracy-irrelevant settings across tested datasets. All Each represents the aggregated distribution under the All Identical Answers setting.
SeViLA [^5] SeViLA uniquely becomes distracted when an empty option is added, suggesting it may have been trained with fixed option counts. Notably, it shows reduced accuracy for all-correct frames in STAR, which aligns with prior findings Goyal et al. (2016); Lin et al. (2023b), where ”blind” language-only models were able to answer image-based questions. demonstrates greater robustness on NExT-QA, maintaining about uniform 20% distribution under harsh attacks. It only becomes unstable when all correct answers are identical (favoring a0) and with empty answers (favoring the last option). Figure 4 shows slightly higher saturation for these conditions. Unlike Video-LLaMA and Video-LLaVA, which rely on multinomial probabilities, SeViLA uses the argmax function, always selecting the first option if probabilities are equal. This explains why it often favors the first option, a0, in identical-answer scenarios.
@@ 4 Methodology
Our bias calibration approach is grounded in established guidelines for developing and validating multiple-choice test items Haladyna (2004), which emphasize balanced, independent, and logically plausible options. This way we ensure that the datasets isolate positional bias and allow our calibration method to be applied effectively. Specifically, we require the following conditions to hold in the MCQA benchmark:
1. Balanced options: All answer options are presented equally without any positional or presentation bias. This is crucial and straightforward to achieve via minimal preprocessing.
2. Independent options: Each option for a given question is independent of others. There should be no overlapping content or logical dependencies among options that influence selection.
3. No logically better or worse options except the correct one: All distractors should be plausible and not trivially inferior or superior to the correct answer. This ensures that the model must rely on genuine reasoning.